# Wake-Informed 3D Path Planning for Autonomous Underwater Vehicles Using A* and Neural Network Approximations

### v1

@@ Abstract
Autonomous Underwater Vehicles (AUVs) encounter significant energy, control and navigation challenges in complex underwater environments, particularly during close-proximity operations, such as launch and recovery (LAR), where fluid interactions and wake effects present additional navigational and energy challenges. Traditional path planning methods fail to incorporate these detailed wake structures, resulting in increased energy consumption, reduced control stability, and heightened safety risks. This paper presents a novel wake-informed, 3D path planning approach that fully integrates localized wake effects and global currents into the planning algorithm. Two variants of the A* algorithm—a current-informed planner and a wake-informed planner— are created to assess its validity and two neural network models are then trained to approximate these planners for real-time applications. Both the A* planners and NN models are evaluated using important metrics such as energy expenditure, path length, and encounters with high-velocity and turbulent regions. The results demonstrate a wake-informed A* planner consistently achieves the lowest energy expenditure and minimizes encounters with high-velocity regions, reducing energy consumption by up to 11.3%. The neural network models are observed to offer computational speedup of 6 orders of magnitude, but exhibit 4.51–19.79% higher energy expenditures and 9.81–24.38% less optimal paths. These findings underscore the importance of incorporating detailed wake structures into traditional path planning algorithms and the benefits of neural network approximations to enhance energy efficiency and operational safety for AUVs in complex 3D domains.
@@
Path Planning , Hydrodynamics , Efficiency , Autonomy , Optimization , Machine Learning , Neural Networks
@@
[label2]organization=Maritime Engineering and Ship Science, Faculty of Engineering and Physical Sciences, University of Southampton, city=Southampton, country=United Kingdom
@@ 1 Introduction
Autonomous underwater vehicles (AUVs) operate in complex, rapidly changing environments, with limited sensory data available. This lack of sensory perception in a complex environment complicates the effective control of these vehicles from a guidance and navigation perspective. This is further compounded by the addition of multiple vehicles or the undertaking of challenging manoeuvres, such as underwater launch and recovery (LAR) of smaller platforms in the unsteady flow wake of a larger platform [1, 2]. Whilst underwater and in operation, AUV platforms are free to move with 6 degrees of freedom and in doing so need to maintain effective surge, heave and sway control, in addition to pitch, yaw and roll [3].
@@
The underwater environment is further complicated by communications challenges, current and dynamic flow effects and the size of the working environment [3]. To ensure safe and efficient navigational paths in the underwater domain, a variety of different path planning methods have been developed or adapted [4]. The majority of these methods seek to optimize travel time [5], energy consumption [6], and obstacle avoidance [7] and can be broadly classified into node-based, sampling-based, or bioinspired, depending on their operational principles and original inspiration.
@@
Limited research has addressed path-planning where vehicle wake structures are included, specifically in 3D applications, opting to instead focus on global current effects [8], geometric obstacle avoidance or simplified [9] flow field structures. The inclusion of flow structures is primarily limited to 2D [8, 9]. Works that have sought to investigate path planning within 3D flows for autonomous applications have stopped short of incorporating the 3D field directly within the model [10], favoring the recommendation of navigational best practices instead. Other research focusing on unmanned aerial vehicle (UAV) navigation in turbulent 3D domains, using the A* algorithm, opted to completely avoid turbulent regions of the path to ensure flight safety [11].
@@
However, for close-proximity underwater operations between vehicles, such as LAR, the turbulent structures are large, strong, and hard to positionally approximate [1, 2]. Due to this, when conducting close-proximity operations between vehicles complete avoidance as detailed in other works [11], is not possible. Subsequently, this necessitates an intelligent, rapid response, and computationally efficient means of navigating near or through wake structures.
In response to these challenges, this paper proposes, develops, and tests a wake-informed, 3D path planning framework for AUVs that integrates detailed hydrodynamic data directly into the planning algorithm. This approach is demonstrated to enable more energy-efficient and safer navigation through complex wake structures. An investigation is then undertaken into the ability of neural networks to approximate the calculated planning approach - showing good initial success. This research highlights the benefits of incorporating full 3D flow information into traditional and novel planning approaches.
@@
To contextualize this work, Section 2 reviews the current state of path planning in the underwater domain and identifies key gaps that motivate the study. Section 3 introduces the computational domain and the specific AUV configurations considered. In Section 4, the development of the wake- and energy-informed planning algorithm is presented, while Section 5 discusses the neural network approach used to approximate the planner for real-time applications. The performance of both the A* planners and their neural network approximations is evaluated in Section 6, followed by a detailed discussion of the findings in Section 7. Finally, Section 8 concludes the paper and outlines potential directions for future research.
@@ 2 Background
A substantial body of literature exists on AUV path planning in complex underwater environments. Many different variations of planning algorithms have been proposed, each seeking to best address a specific challenge that impacts underwater vehicles. The vast majority of these works can be first attributed to aerial and terrestrial path planning algorithms before modification for underwater applications. These approaches can be further divided according to their method of operation, such as node-based, sampling-based, and bioinspired planning [12].
2.1 Path Planning Methods
Node-based methods, such as the A* algorithm, are widely used due to their ability to find optimal paths in discretized environments [13]. Node based methods typically operate on a grid or graph representation of the environment and consider factors such as travel time, energy consumption and obstacle avoidance. Sampling-based methods, like Rapidly-exploring Random Trees (RRT) and Probabilistic Roadmaps (PRM), are particularly useful in high-dimensional spaces and can handle complex constraints [14, 15]. The sampling-based methods are effective in real-time applications due to their probabilistic completeness and ability to find feasible paths quickly, which some computationally intensive methods, such as bioinspired methods, can struggle with. The bioinspired algorithms, such Ant Colony Optimization and Genetic Algorithms, mimic natural processes to find efficient paths [16, 17]. These methods are advantageous in handling dynamic environments and optimizing multiple objectives simultaneously, but are not guaranteed to provide the optimal path in all cases.
2.2 Incorporating Environmental Effects
Within each path planner, multiple effects can be modeled or accounted for. For the operation of underwater vehicles, environmental effects are important. The environment that an AUV exists in can affect its ability to safely or effectively navigate. Previous works have incorporated a range of effects, seeking to produce an informed planner. These have included the inclusion of ocean currents to help identify optimal time and energy optimal routes [18, 4, 19]. However, the dynamic current effects are often incorporated using simplified or 2D representations of flow fields. Garau et al. [13] presented an A* path planning algorithm that considers spatially variable current fields, but this field was limited to 2D, neglecting the 3D complexity of real flow structures. Similarly, Petres et al. [12] proposed a method that incorporates ocean currents into path planning but relies on simplified 2D current models.
Some research has extended the 2D effects in path planning to 3D. This has primarily been to account for depth variations and 3D obstacles [20, 21]. Zadeh et al. [20] developed a 3D path planning method based on ocean currents and differential evolution algorithms but did not fully incorporate complex hydrodynamic effects directly. Zeng et al. [22] sought to extend 2D current data to three dimensions, undertaking this by simplified linear approximations of the 2D current-only flow surface, extended into the depth axis. Li and Yu [21] proposed a significantly more detailed approach that included current structures, travel distance and object constraints in a 3D underwater environment with ocean currents, yet it again assumed a simplified current flow field and did not include wake structures from the objects or vehicles in the field.
2.3 Limitations in Current Research
For underwater applications, and specifically those with two or more interacting vehicles, an understanding of the best path through a fully realized 3D field is important. These operations are common and include, but are not limited to launch and recovery, formation flying and multi-vehicle cable/pipe inspection underwater. In these applications, the 3D flow field, wake structures and shed hydrodynamics from vehicles may pose control, safety and energy efficiency effects that are important to account for [23].
To date, limited research has sought to account for these effects within a path planner that incorporates full 3D hydrodynamic models, including the wake structures of nearby vessels. Existing methods either simplify the flow field in relation to its dimensionality (3D to 2D), include only the current, avoiding turbulent regions altogether, or use approximated hydrodynamic models [24, 25]. These simplifications may not accurately account for dynamics of importance in close proximity operations, such as berthing. Research focusing on UAV navigation in turbulent 3D aerial domains using the A* algorithm opted to completely avoid turbulent regions to ensure flight safety [26]. When conducting close proximity operations between vehicles, such as docking or formation keeping, avoiding these regions is not always a practical option as it can limit the possible safe approach trajectories that the vehicle can employ.
Building on these gaps, it is theorized that a machine learning-based planner, trained on trajectories generated by a high-fidelity A* planner, could learn to replicate or closely approximate the performance of optimal solutions. Such an approach would enable the rapid prediction of feasible, energy-efficient paths without the need for computationally expensive online searches, potentially bringing real-time 3D wake-informed navigation within reach.
2.4 Contribution
This paper makes three primary contributions, namely it:
  1. 1.
    proposes a wake-informed A* path planning algorithm that integrates full 3D hydrodynamic data, including complex wake structures, to achieve more energy-efficient and safer AUV navigation. Bridging the gap between simplified flow models and the complex realities of 3D underwater environments is crucial for operations involving significant wake effects [1, 2].
  2. 2.
    introduces a machine learning framework trained on wake- and current-informed A* solutions, enabling rapid trajectory generation that approximates near-optimal paths in real-time, levering existing techniques for ML informed planning [27, 28].
  3. 3.
    develops a robust set of metrics to holistically evaluate energy consumption, path length, turbulence exposure, and computational efficiency, advancing the understanding of AUV path planning performance in complex 3D flows.
@@ 3 Computational Domain
This method of recovery has already been proposed and hydrodynamically assessed [1, 2], but remains challenging due to the fluid structures present during the manoeuvre which are excluded in current path planners. The propeller race and vehicle wake of the lead vessel presents a complex navigational challenge to the approaching vehicle. Approaching with limited knowledge of the strong flow structure, instead accounting only for the current, can complicate attempts to make a safe recovery. Thus, a smart means of avoidance is required. This presents an optimal task by which to assess, benchmark and evaluate efficient methods for wake-informed path planning in 3D domains.
@@
The LAR domain is modeled as a pseudo-static environment. The larger vehicle, an extra-large uncrewed underwater vehicle (XLUUV) is located within the center of the computational grid. This grid represents a 155 $\times$ 155 $\times$ by 155 meter physical domain. The larger vehicle to which the recovery is made is modeled to be 22 meters in length, 2.2 meters wide and 2.7 meters tall. This vehicle is equipped with an INSEAN E1619 propeller, based on the work of [1]. It is modeled with a central payload bay that is 5.5 meters long, 1.5 meters wide and 2.2 meters tall. It is assumed that for each planned trajectory, the XLUUV velocity and heading are constant. These manoeuvres take place at a modelled depth of 100m, with a constant current and no wave or free surface effects.
@@
This payload bay is accessible from the bottom, mimicking the payload bay LAR as assessed by [1] and [2], which indicated such a maneuver is possible. This domain is contained within a dataset of 500 distinct flow conditions, as discussed in Section 4.2. The velocity magnitude contained within the ANSYS Fluent CFD data has been interpolated from the mesh nodes of each simulation onto the regularly spaced 3D grid structure used in the planner. This was undertaken using a nearest-neighbor approach between vertexes of the CFD data and the 3D grid structure as discussed in Section 4.1. Example trajectories through the domain are provided in Fig. 1 and 2 which show the larger vehicle, the wake structure and an example approach trajectory.
Figure 1: Front side view of the domain. Larger vehicle geometry indicated in blue, contained in the wake structure of the body and propeller (green). Indicated in red is an approach trajectory of the vehicle attempting recovery.
Figure 2: Bottom view of the approach. Larger vehicle geometry indicated in blue, contained in the wake structure of the body and propeller (green). Indicated in red is an approach trajectory of the vehicle attempting recovery. Payload bay location onboard the larger vehicle necessitates a transition through the boundary layer, whilst avoidance of the far field wake is ideal

### v2

@@ Abstract
Autonomous Underwater Vehicles (AUVs) encounter significant energy, control and navigation challenges in complex underwater environments, particularly during close-proximity operations, such as launch and recovery (LAR), where fluid interactions and wake effects present additional navigational and energy challenges. Traditional path planning methods fail to incorporate these detailed wake structures, resulting in increased energy consumption, reduced control stability, and heightened safety risks. This paper presents a novel wake-informed, 3D path planning approach that fully integrates localized wake effects and global currents into the planning algorithm. Two variants of the A* algorithm – a current-informed planner and a wake-informed planner – are created to assess its validity and two separate neural network models are then trained, each designed to approximate one of the A* planner variants (current-informed and wake-informed respectively), enabling potential real time-application. Both the A* planners and NN models are evaluated using important metrics such as energy expenditure, path length, and encounters with high-velocity and turbulent regions. The results demonstrate a wake-informed A* planner consistently achieves the lowest energy expenditure and minimizes encounters with high-velocity regions, reducing energy consumption by up to 11.3%. The neural network models are observed to offer computational speedup of 6 orders of magnitude, but exhibit 4.51–19.79% higher energy expenditures and 9.81–24.38% less optimal paths. These findings underscore the importance of incorporating detailed wake structures into traditional path planning algorithms and the benefits of neural network approximations to enhance energy efficiency and operational safety for AUVs in complex 3D domains.
@@
Path Planning , Hydrodynamics , Autonomous Underwater Vehicles , Optimization , Machine Learning
@@
[label2]organization=Maritime Engineering, Faculty of Engineering and Physical Sciences, University of Southampton, city=Southampton, country=United Kingdom
@@ 1 Introduction
Autonomous underwater vehicles (AUVs) operate in complex, rapidly changing environments, with limited sensory data available. This lack of sensory perception in a complex environment complicates the effective control of these vehicles from a guidance and navigation perspective. This is further compounded by the addition of multiple vehicles or the undertaking of challenging manoeuvres, such as underwater launch and recovery (LAR) of smaller platforms in the unsteady flow wake of a larger platform (Cooper-Baldock et al., 2023, 2024). Whilst underwater and in operation, AUV platforms are free to move with 6 degrees of freedom and in doing so need to maintain effective surge, heave and sway control, in addition to pitch, yaw and roll (Petres et al., 2007a).
@@
The underwater environment is further complicated by communications challenges, current and dynamic flow effects and the size of the working environment (Petres et al., 2007a). To ensure safe and efficient navigational paths in the underwater domain, a variety of different path planning methods have been developed or adapted (Alvarez et al., 2004a). The majority of these methods seek to optimize travel time (Latombe, 1991), energy consumption (Bellingham and Willcox, 1996), and obstacle avoidance (Hernández et al., 2015) and can be broadly classified into node-based, sampling-based, or bio-inspired, depending on their operational principles and original inspiration.
@@
Limited research has addressed path-planning where vehicle wake structures are included, specifically in 3D applications, opting to instead focus on global current effects (Lermusiaux et al., 2016), geometric obstacle avoidance or simplified (Qi et al., 2022) flow field structures. The inclusion of flow structures is primarily limited to 2D (Lermusiaux et al., 2016; Qi et al., 2022). Works that have sought to investigate path planning within 3D flows for autonomous applications have stopped short of incorporating the 3D field directly within the model (Wang et al., 2023), favoring the recommendation of navigational best practices instead. Other research focusing on unmanned aerial vehicle (UAV) navigation in turbulent 3D domains, using the A* algorithm, opted to completely avoid turbulent regions of the path to ensure flight safety (Pensado et al., 2024a).
@@
However, for close-proximity underwater operations between vehicles, such as LAR, the turbulent structures are large, strong, and hard to positionally approximate (Cooper-Baldock et al., 2023, 2024). Due to this, when conducting close-proximity operations between vehicles, complete avoidance as used in other works (Pensado et al., 2024a) is not possible. Subsequently, this necessitates an intelligent, rapid, and computationally efficient means of navigating near or through wake structures borne from other vessels.
1.1 Limitations in Current Research
For underwater applications, and specifically those with two or more interacting vehicles, an understanding of the best path through a fully realized 3D field is important. These operations are common and include, but are not limited to, launch and recovery, formation flying and multi-vehicle underwater cable and pipe inspection operations. In these applications, the 3D flow field, including wake structures and shed hydrodynamics each of the vehicles may pose control, safety and energy efficiency effects that are important to account for (Bhattacharyya, 1978).
To date, limited research has sought to model these effects within a path planner, at a level sufficient to encapsulate the complex 3D hydrodynamic physics, including the wake structures of nearby vessels. Existing methods either simplify the flow field in relation to its dimensionality (3D to 2D (Garau et al., 2005; Petres et al., 2007b)), include only the current, avoiding turbulent regions altogether (Pensado et al., 2024b), or use approximated hydrodynamic models (Zhang et al., 2022; Ronghua et al., 2024; Zadeh et al., 2018; Zeng et al., 2018; Li and Yu, 2023). Research focusing on UAV navigation in turbulent 3D aerial domains using the A* algorithm opted to completely avoid turbulent regions entirely to ensure flight safety (Pensado et al., 2024b). However, when conducting close proximity operations between vehicles, such as docking, formation flying or collaborative tasks, avoiding these regions is not always a practical or possible option. Vehicles may be required to navigate in close proximity to achieve their goals. These existing models may be inadequate to capture the fluid dynamic impacts present in multi-vehicle operations, where strong 3D wake interactions can significantly impact AUV control, safety, and energy consumption (Cooper-Baldock et al., 2023, 2024; Bhattacharyya, 1978).
To investigate whether a detailed knowledge of these wake hydrodynamics is required, and the benefits of doing so, two A* path planners will be developed for comparison – one in which only the global current is considered, and one where the detailed 3D wake structure is accounted for. Additionally, the use of machine learning based planners, trained on the A* trajectories, will also be investigated. Such an approach may enable the rapid prediction of feasible, energy-efficient paths without the need for computationally expensive A* searches, potentially bringing real-time 3D wake-informed navigation within reach.
1.2 Contributions
To address these scientific gaps, this paper makes the following primary contributions:
  1. 1.
    Proposes a wake-informed A* path planning algorithm that integrates full 3D hydrodynamic data, including complex wake structures, to achieve more energy-efficient and safer AUV navigation. Bridging the gap between simplified flow models and the complex realities of 3D underwater environments is crucial for operations involving significant wake effects (Cooper-Baldock et al., 2023, 2024).
  2. 2.
    Introduces a machine learning framework trained on wake- and current-informed A* trajectories, enabling rapid real time path generation that approximates the near-optimal paths in real-time, levering existing techniques for ML informed planning (Tai et al., 2017; Pfeiffer et al., 2017).
  3. 3.
    Develops a robust set of metrics to holistically evaluate energy consumption, path length, turbulence exposure, and computational efficiency, advancing the understanding of AUV path planning performance in complex 3D domains.
@@ 1.2 Contributions
To contextualize this work, Section 2 reviews the current state of path planning in the underwater domain and identifies key gaps that motivate the study. Section 3 introduces the computational domain and the specific AUV configurations considered. In Section 4, the development of the wake- and energy-informed planning algorithm is presented, while Section 5 discusses the neural network approach used to approximate the planners for real-time applications. The performance of both the A* planners and their neural network counterparts is evaluated in Section 6, followed by a detailed discussion of the findings in Section 7. Finally, Section 8 concludes the paper and outlines potential directions for future research.
@@ 2 Background
Effective path planning is crucial for Autonomous Underwater Vehicle (AUV) operations in complex environments, aiming to optimize metrics like travel time, energy consumption, and safety (Latombe, 1991; Bellingham and Willcox, 1996; Hernández et al., 2015). Table 1 summarizes relevant path planning literature, highlighting the methods, considered environmental effects, dimensionality, and key limitations, particularly concerning the integration of detailed 3D hydrodynamics. Common approaches adapt terrestrial or aerial methods, broadly categorized as node-based (e.g. A* (Garau et al., 2005)), sampling-based (e.g. RRT (LaValle, 2006; Janoš et al., 2021)), or bio-inspired methods (e.g. Ant Colony Optimization (Akka and Khaber, 2018; Alvarez et al., 2004b)). The domain that AUVs operate in is complex, as they are affected by global current effects(Kruger et al., 2007), wake structures (Cooper-Baldock et al., 2024), local obstacles that must be avoided and the hydrodynamics of other vessels (Cooper-Baldock et al., 2024).
While numerous studies have integrated ocean currents to find time or energy-optimal routes (Kruger et al., 2007; Alvarez et al., 2004a; Kulkarni and Lermusiaux, 2020), these often rely on simplified 2D representations (Garau et al., 2005; Petres et al., 2007b) or extend currents to 3D using approximations that neglect the full dimensionality of the flow, omitting complex, localized hydrodynamic features like vehicle wakes (Zadeh et al., 2018; Zeng et al., 2018; Li and Yu, 2023). Research involving turbulent aerial domains has been observed to avoid the turbulent regions entirely (Pensado et al., 2024b), an approach that is impractical or impossible for close-proximity underwater maneuvers such as launch and recovery (Cooper-Baldock et al., 2023, 2024). More recently, machine learning techniques and specifically neural networks (NNs), have been applied to path planning tasks to capitalizing on computational speed and hardware requirement reductions (Tai et al., 2017; Pfeiffer et al., 2017). However, these are frequently applied to simplified 2D grids, and have a focus on obstacle avoidance rather than navigation of complex 3D flow fields.
Table 1: Summary of Relevant Path Planning Literature for Autonomous Maritime Vehicles.
Reference(s) Method Included Effects Dim. Pros Limitations
(Garau et al., 2005) A* (Node-based) Spatially variable currents 2D Considers current effects on path cost. Limited to 2D; No wake structures.
(Petres et al., 2007b) Review (Multiple) Ocean currents (General focus) 2D Reviews various methods for current planning. Highlights reliance on simplified 2D models.
(Alvarez et al., 2004a) Evolutionary Ocean currents 2D Optimize time/energy with currents. Limited to 2D; No wake structures.
(Kruger et al., 2007) A* (Node-based) Estuary currents 3D Considers relative motion and energy Limited 3D flow model; No wakes
(Kulkarni and Lermusiaux, 2020) Optimization Ocean currents 3D Optimize time/energy with currents. Uses simplified current fields; No detailed wakes.
(Zadeh et al., 2018) Differential Evolution Ocean currents, 3D obstacles 3D Handles 3D obstacles and currents. Did not fully integrate complex hydrodynamic effects.
(Zeng et al., 2018) A* Extended 2D current data 3D Attempted 3D current representation. Simplified linear extension of 2D data; No wake.
(Li and Yu, 2023) Optimization Currents, Obstacles, Travel Dist. 3D Detailed 3D constraints considered. Assumed simplified current flow; No wake structures.
(Pensado et al., 2024b) A* Turbulence (Aerial UAV focus) 3D Addresses navigation in turbulence. Avoids turbulent regions entirely; Not wake-specific.
(Tai et al., 2017; Pfeiffer et al., 2017) Neural Networks (NN) Obstacles, Environment Features 2D/3D Fast computation; Learning-based. Often applied to simpler env./grids; Not focused on detailed 3D wakes.
This Work A* & NN Currents & Detailed Wake 3D Integrates full 3D wake data; Compares A*/NN. Focuses on specific LAR scenario
The literature summarized in Table 1 indicates that while progress has been made in incorporating environmental factors like currents and navigating 3D spaces, a critical gap remains in path planning methodologies that fully integrate detailed, localized 3D hydrodynamic effects, specifically the complex wake structures generated by nearby vehicles. These wakes significantly impact AUV control, safety, and energy efficiency during close-proximity operations (Bhattacharyya, 1978; Cooper-Baldock et al., 2023, 2024), yet are typically simplified or ignored. This research directly addresses this limitation by developing and evaluating path planners—both high-fidelity A* algorithms and computationally efficient neural network approximations—that explicitly incorporate detailed 3D wake data alongside global currents to improve AUV navigation performance in such challenging scenarios.
@@ 3 Computational Domain
This method of recovery has already been proposed and hydrodynamically assessed (Cooper-Baldock et al., 2023, 2024), but remains challenging due to the fluid structures present during the manoeuvre which are excluded in current path planners. The propeller race and vehicle wake of the lead vessel presents a complex navigational challenge to the approaching vehicle. Approaching with limited knowledge of the strong flow structure, instead accounting only for the current, can complicate attempts to make a safe recovery. Thus, a smart means of avoidance is required. This presents an optimal task by which to assess, benchmark and evaluate efficient methods for wake-informed path planning in 3D domains.
@@
The LAR domain is modeled as a pseudo-static environment. The larger vehicle, an extra-large uncrewed underwater vehicle (XLUUV) is located within the center of the computational grid. This grid represents a 155 $\times$ 155 $\times$ by 155 meter physical domain. The larger vehicle to which the recovery is made is modeled to be 22 meters in length, 2.2 meters wide and 2.7 meters tall. This vehicle is equipped with an INSEAN E1619 propeller, based on the work of (Cooper-Baldock et al., 2024). It is modeled with a central payload bay that is 5.5 meters long, 1.5 meters wide and 2.2 meters tall. It is assumed that for each planned trajectory, the XLUUV velocity and heading are constant. These manoeuvres take place at a modeled depth of 100m, with a constant current and no wave or free surface effects.
@@
This payload bay is accessible from the bottom, mimicking the payload bay LAR as assessed by (Cooper-Baldock et al., 2024) and (Cooper-Baldock et al., 2023), which indicated such a maneuver is possible. This domain is contained within a dataset of 500 distinct flow conditions, as discussed in Section 4.3. The velocity magnitude contained within the ANSYS Fluent CFD data has been interpolated from the mesh nodes of each simulation onto the regularly spaced 3D grid structure used in the planner. This was undertaken using a nearest-neighbor approach between vertexes of the CFD data and the 3D grid structure as discussed in Section 4.1. Example trajectories through the domain are provided in Fig. 1 and 2 which show the larger vehicle, the wake structure and an example approach trajectory.
Figure 1: Side view of the optimal A* wake-informed path planner trajectory for a given flow state. The AUV follows the trajectory indicated in red, minimizing its interaction with the wake structure of the XLUUV, where the high speed flow is indicated in green. The XLUUV body geometry is visible in blue, through the transparent flow field in green. In this case, the AUV must navigate to the goal location within the XLUUVs’ central payload bay, necessitating a transition through the boundary layer.
Figure 2: Bottom up view of the optimal A* wake-informed path planner trajectory for a given flow state. The AUV follows the trajectory indicated in red, minimizing its interaction with the wake structure of the XLUUV, where the high speed flow is indicated in green. The XLUUV body geometry is visible in blue, through the transparent flow field in green. In this case, the AUV must navigate to the goal location within the XLUUVs’ central payload bay, necessitating a transition through the boundary layer.